# Incorporating Arbitrary Matrix Group Equivariance into KANs

### v3

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have seen great success in scientific domains thanks to spline activation functions, becoming an alternative to Multi-Layer Perceptrons (MLPs). However, spline functions may not respect symmetry in tasks, which is crucial prior knowledge in machine learning. In this paper, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), a method for incorporating arbitrary matrix group equivariance into KANs, aiming to broaden their applicability to more fields. We first construct gated spline basis functions, which form the EKAN layer together with equivariant linear weights, and then define a lift layer to align the input space of EKAN with the feature space of the dataset, thereby building the entire EKAN architecture. Compared with baseline models, EKAN achieves higher accuracy with smaller datasets or fewer parameters on symmetry-related tasks, such as particle scattering and the three-body problem, often reducing test MSE by several orders of magnitude. Even in non-symbolic formula scenarios, such as top quark tagging with three jet constituents, EKAN achieves comparable results with state-of-the-art equivariant architectures using fewer than $40\%$ of the parameters, while KANs do not outperform MLPs as expected.
Equivariant neural network, Kolmogorov-Arnold network, Lie theory
@@ 1 Introduction
Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024b, a) are a novel type of neural network inspired by the Kolmogorov-Arnold representation theorem (Tikhomirov, 1991; Braun & Griebel, 2009), which offers an alternative to Multi-Layer Perceptrons (MLPs) (Haykin, 1998; Cybenko, 1989; Hornik et al., 1989). Unlike MLPs, which utilize fixed activation functions on nodes, KANs employ learnable activation functions on edges, replacing the linear weight parameters entirely with univariate functions parameterized as splines (De Boor & De Boor, 1978). On the other hand, each layer of KANs can be viewed as spline basis functions followed by a linear layer (Dhiman, 2024). This architecture allows KANs to achieve better accuracy in symbolic formula representation tasks compared with MLPs, particularly in function fitting and scientific applications. Subsequent works based on KANs have demonstrated superior performance in other areas, such as sequential data (Vaca-Rubio et al., 2024; Genet & Inzirillo, 2024b, a; Xu et al., 2024), graph data (Bresson et al., 2024; De Carlo et al., 2024; Kiamari et al., 2024; Zhang & Zhang, 2024), image data (Cheon, 2024b, a; Azam & Akhtar, 2024; Li et al., 2024a; Seydi, 2024; Bodner et al., 2024), and so on.
However, KANs themselves perform poorly on non-symbolic formula representation tasks (Yu et al., 2024). One reason for this is that splines struggle to respect data type and symmetry, both of which play important roles in machine learning. Many recent works utilize symmetry in data to design network architectures, achieving better efficiency and generalization on specific tasks. For example, Convolutional Neural Networks (CNNs) (LeCun et al., 1989) and Group equivariant Convolutional Neural Networks (GCNNs) (Cohen & Welling, 2016a) leverage translational and rotational symmetries in image data, while DeepSets (Zaheer et al., 2017) and equivariant graph networks (Maron et al., 2018; Keriven & Peyré, 2019; Satorras et al., 2021) exploit the permutation symmetry in set and graph data. Equivariant Multi-Layer Perceptrons (EMLP) (Finzi et al., 2021) propose a general method that allows MLPs to be equivariant with respect to arbitrary matrix groups for specific data types, thereby unifying the aforementioned specialized network architectures.
@@ 2 Background
where $\oplus$ is the direct sum, $\otimes$ is the Kronecker product, and $\boxplus$ is the Kronecker sum.
@@ 3 Related work
Equivariant networks have gained significant attention in recent years due to their ability to respect and leverage symmetries in data. GCNNs (Cohen & Welling, 2016a) embed discrete group equivariance into traditional CNNs through group convolutions. Steerable CNNs (Cohen & Welling, 2016b) introduce steerable filters, which provide a more flexible and efficient way to achieve equivariance compared with GCNNs. Subsequently, SFCNNs (Weiler et al., 2018b) and E(2)-equivariant steerable CNNs (Weiler & Cesa, 2019) extend GCNNs and steerable CNNs to continuous group equivariance, while 3D Steerable CNNs (Weiler et al., 2018a) extend these models to 3D volumetric data. On the other hand, some works use partial differential operators (PDOs) to construct equivariant networks (Shen et al., 2020, 2021, 2022; He et al., 2022; Li et al., 2024b). Furthermore, equivariant self-supervised learning (Wang et al., 2020b; Dangovski et al., 2021; Lee et al., 2022; Garrido et al., 2023; Gupta et al., 2023) has also achieved outstanding results. Based on these theoretical frameworks, equivariant networks are widely applied in various fields, such as mathematics (Zhao et al., 2022), physics (Wang et al., 2020a), biochemistry (Bekkers et al., 2018; Winkels & Cohen, 2019; Graham et al., 2020), and others. EMLP (Finzi et al., 2021) embeds matrix group equivariance into MLPs layerwise, which we discuss in detail in Appendix B.
@@ 4.1 Space structures
where $p_{i,a},q_{i,a},p_{o,a},q_{o,a},c_{i},c_{o},A_{i},A_{o}\in\mathbb{N}$, $p_{i,a}^{2}+q_{i,a}^{2}>0$, $p_{o,a}^{2}+q_{o,a}^{2}>0$, and $cT_{0}=\underbrace{T_{0}\oplus T_{0}\oplus\dots\oplus T_{0}}_{c}$.
We have to emphasize that the actual input/output feature does not lie within $U_{i}$/$U_{o}$. To align with gated basis functions, we add a gate scalar $T_{0}$ to each non-scalar term $T(p_{i,a},q_{i,a})$/$T(p_{o,a},q_{o,a})$ in $U_{i}$/$U_{o}$ to obtain the actual input/output space. The mechanism behind this approach will be discussed in detail in Section 4.2. We denote this actual input/output space as the gated input/output space $U_{gi}$/$U_{go}$:
@@ 4.2 Gated basis functions
The following theorem guarantees the equivariance between the gated input space and the post-activation space (see Appendix C for proof).
@@ 4.3 Equivariant linear weights
where $\rho_{go,i}=\rho_{go}\otimes\rho_{i}^{*}$ is the group representation of $U_{go}\otimes U_{i}^{*}$. We use the same method as EMLP (Finzi et al., 2021) to solve for the equivariant basis $Q$ and equivariant projector $P=QQ^{\top}$ of $\mathrm{vec}(W_{b})$. Similar to the transition from Equation 27 to Equation 28 (in Appendix B), we decompose the group representation $\rho_{go,i}(g)$ into discrete and infinitesimal generators to obtain the following constraint:
@@
We summarize the architecture of the EKAN layer in Figure 2 (Right). In this example, the EKAN layer is equivariant with respect to a 2-dimensional matrix group $\widetilde{G}$ (such as the $\mathrm{SO}(2)$ group). The user specifies the input space $U_{i}=T_{0}\oplus T_{1}$ (where we abbreviate $T(p,0)=V^{p}$ as $T_{p}$), which represents a scalar space and a vector space, and specifies the output space $U_{o}=T_{2}$, which represents a matrix space. Then, the gated input space $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ and the gated output space $U_{go}=T_{2}\oplus T_{0}$ each add a gate scalar to the vector space $T_{1}$ and the matrix space $T_{2}$. The basis functions are applied to the gate scalars of $T_{0}$ (itself) and $T_{1}$, which are then multiplied by the original terms to obtain the post-activation space $U_{m}=3U_{i}$. The linear weights $W\in\mathbb{R}^{5\times 9}$ between $U_{m}$ and $U_{go}$ are within the subspace determined by Section 4.3 to ensure equivariance. Similar to KANs (Liu et al., 2024b), EKAN updates grids based on the input activations, which we discuss in detail in Appendix D.
@@ 5 EKAN architecture
However, the actual input/output features of the EKAN layer stack lie in $U_{gi}$/$U_{go}$. Therefore, we need to add extensions to align the network with the dataset. First, the gate scalars of the actual output feature are directly dropped to obtain the final output label of EKAN, which resides in $U_{o}$. Then, we add a lift layer before the first layer to preprocess the raw input feature of EKAN, which is essentially an equivariant linear layer between $U_{i}$ and $U_{gi}$ (see Section 4.3).
@@ 6 Experiments
In this section, we evaluate the performance of EKAN on regression and classification tasks with known symmetries. Compared with general models such as MLPs, KANs, and equivariant architectures like EMLP (Finzi et al., 2021) and CGENN (Ruhe et al., 2023), EKAN achieves lower test loss and higher test accuracy with smaller datasets or fewer parameters.
@@ 6.1 Particle scattering
In electron-muon scattering, we can observe the four-momenta of the incoming electron, incoming muon, outgoing electron, and outgoing muon, denoted as $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\in\mathbb{R}^{4}$, respectively. We aim to predict the matrix element, which is proportional to the cross-section (Finzi et al., 2021): $|\mathcal{M}|^{2}\propto[p^{\mu}\tilde{p}^{\nu}-(p^{\alpha}\tilde{p}_{\alpha}-p^{\alpha}p_{\alpha})g^{\mu\nu}][q_{\mu}\tilde{q}_{\nu}-(q^{\alpha}\tilde{q}_{\alpha}-q^{\alpha}q_{\alpha})g_{\mu\nu}]$. According to Einstein’s summation convention, in a monomial, if an index appears once as a superscript and once as a subscript, it indicates summation over that index. The metric tensor is given by $g_{\mu\nu}=g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, and $a_{\mu}=g_{\mu\nu}a^{\nu}=(a^{0},-a^{1},-a^{2},-a^{3})$. The matrix element is invariant under Lorentz transformations. In other words, this task exhibits $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=4T_{1}$ and the label space $U_{o}=T_{0}$.
We embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN. Models are evaluated on synthetic datasets with different training set sizes, which are generated by sampling $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\sim\mathcal{N}(0,\frac{1}{4^{2}})$. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. Although the lift layer increases the parameter overhead, we set the width of the middle layer in EKAN to $n_{1}=1000$ (shape as $[16,1000,1]$, and the software will automatically calculate the appropriate feature space structure based on the user-specified dimension), and set the width of the middle layer in KAN to $n_{1}=3840$ (shape as $[16,3840,1]$) to keep the parameter count similar. Both EMLP and MLP have the depth of $L=4$ and the middle layer width of $n_{1}=n_{2}=n_{3}=384$ (shape as $[16,384,384,384,1]$). In these settings, EKAN ($435$k) has fewer parameters than EMLP ($450$k) and KAN ($461$k). We provide more implementation details in Section E.1.
@@ 6.2 Three-body problem
[…]
@@
EMLP-$\mathrm{SO}^{+}(1,3)$ $65.48\pm 1.21$ $72.59\pm 0.84$ $74.40\pm 0.26$ $76.34\pm 0.14$ $77.10\pm 0.02$ $133$K
EMLP-$\mathrm{SO}(1,3)$ $61.86\pm 5.92$ $73.09\pm 0.92$ $74.37\pm 0.17$ $76.46\pm 0.12$ $\mathbf{77.12\pm 0.04}$
EMLP-$\mathrm{O}(1,3)$ $62.66\pm 7.35$ $73.65\pm 1.01$ $74.22\pm 0.53$ $76.26\pm 0.05$ $\mathbf{77.12\pm 0.04}$
KAN $49.89\pm 0.39$ $49.91\pm 0.43$ $49.89\pm 0.37$ $50.00\pm 0.02$ $49.84\pm 0.25$ $35$K
@@
CGENN $62.63\pm 2.24$ $68.74\pm 0.77$ $70.29\pm 1.29$ $75.10\pm 0.47$ $77.05\pm 0.03$ $85$K
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $\mathbf{71.92\pm 0.88}$ $\mathbf{73.98\pm 0.39}$ $\mathbf{76.15\pm 0.11}$ $\mathbf{76.69\pm 0.08}$ $76.93\pm 0.02$ $34$K
@@
The study of the three-body problem on a plane (Greydanus et al., 2019) focuses on the motion of three particles, with their center of mass at the origin, under the influence of gravity. Their trajectories are chaotic and cannot be described by an analytical solution. Specifically, we observe the motion states of three particles over the past four time steps, denoted as $\{q_{i1},p_{i1},q_{i2},p_{i2},q_{i3},p_{i3}\}_{i=t-4}^{t-1}$, and predict their motion states at time $t$, denoted as $\{q_{t1},p_{t1},q_{t2},p_{t2},q_{t3},p_{t3}\}$. Here, $q_{ij}\in\mathbb{R}^{2}$ and $p_{ij}\in\mathbb{R}^{2}$ indicate the position and momentum coordinates of the $j$-th particle at time $i$, respectively. The dataset contains $30$k training samples and $30$k test samples. When the input motion states are simultaneously rotated by a certain angle or reflected along a specific axis, the output motion states should undergo the same transformation. Therefore, this task has $\mathrm{O}(2)$ equivariance (see Section A.1 for more details), with the feature space $U_{i}=4\times 6T_{1}=24T_{1}$ and the label space $U_{o}=6T_{1}$.
We embed the group $\mathrm{O}(2)$ and its subgroup $\mathrm{SO}(2)$ equivariance into EKAN and EMLP. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$, while both EMLP and MLP have the depth of $L=4$. The number of parameters is controlled by adjusting the middle layer width $N$ for comparison (the shape of EKAN and KAN is $[48,N,12]$, while the shape of EMLP and MLP is $[48,N,N,N,12]$). More implementation details can be found in Section E.2.

### v4

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have seen great success in scientific domains thanks to spline activation functions, becoming an alternative to Multi-Layer Perceptrons (MLPs). However, spline functions may not respect symmetry in tasks, which is crucial prior knowledge in machine learning. In this paper, we propose Equivariant Kolmogorov-Arnold Networks (EKAN), a method for incorporating arbitrary matrix group equivariance into KANs, aiming to broaden their applicability to more fields. We first construct gated spline basis functions, which form the EKAN layer together with equivariant linear weights, and then define a lift layer to align the input space of EKAN with the feature space of the dataset, thereby building the entire EKAN architecture. Compared with baseline models, EKAN achieves higher accuracy with smaller datasets or fewer parameters on symmetry-related tasks, such as particle scattering and the three-body problem, often reducing test MSE by several orders of magnitude. Even in non-symbolic formula scenarios, such as top quark tagging with three jet constituents, EKAN achieves comparable results with state-of-the-art equivariant architectures using fewer than $40\%$ of the parameters, while KANs do not outperform MLPs as expected. Code and data are available at https://github.com/hulx2002/EKAN.
Machine Learning, ICML
@@ 1 Introduction
Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024b, a) are a novel type of neural network inspired by the Kolmogorov-Arnold representation theorem (Tikhomirov, 1991; Braun & Griebel, 2009), which offers an alternative to Multi-Layer Perceptrons (MLPs) (Haykin, 1998; Cybenko, 1989; Hornik et al., 1989). Unlike MLPs, which utilize fixed activation functions on nodes, KANs employ learnable activation functions on edges, replacing the linear weight parameters entirely with univariate functions parameterized as splines (De Boor, 1978). On the other hand, each layer of KANs can be viewed as spline basis functions followed by a linear layer (Dhiman, 2024). This architecture allows KANs to achieve better accuracy in symbolic formula representation tasks compared with MLPs, particularly in function fitting and scientific applications. Subsequent works based on KANs have demonstrated superior performance in other areas, such as sequential data (Vaca-Rubio et al., 2024; Genet & Inzirillo, 2024b, a; Xu et al., 2024), graph data (Bresson et al., 2024; De Carlo et al., 2024; Kiamari et al., 2024; Zhang & Zhang, 2024), image data (Cheon, 2024b, a; Azam & Akhtar, 2024; Li et al., 2024a; Seydi, 2024; Bodner et al., 2024), and so on.
However, KANs themselves perform poorly on non-symbolic formula representation tasks (Yu et al., 2024). One reason for this is that splines struggle to respect data type and symmetry, both of which play important roles in machine learning. Many recent works utilize symmetry in data to design network architectures, achieving better efficiency and generalization on specific tasks. For example, Convolutional Neural Networks (CNNs) (LeCun et al., 1989) and Group equivariant Convolutional Neural Networks (GCNNs) (Cohen & Welling, 2016) leverage translational and rotational symmetries in image data, while DeepSets (Zaheer et al., 2017) and equivariant graph networks (Maron et al., 2019; Keriven & Peyré, 2019; Satorras et al., 2021) exploit the permutation symmetry in set and graph data. Equivariant Multi-Layer Perceptrons (EMLP) (Finzi et al., 2021) propose a general method that allows MLPs to be equivariant with respect to arbitrary matrix groups for specific data types, thereby unifying the aforementioned specialized network architectures.
@@ 2 Background
where $\oplus$ is the direct sum, $\otimes$ is the Kronecker product, and $\boxplus$ is the Kronecker sum. We provide concrete examples of the space structure in Appendix B to help readers understand it intuitively.
@@ 3 Related Works
Equivariant networks have gained significant attention in recent years due to their ability to respect and leverage symmetries in data. GCNNs (Cohen & Welling, 2016) embed discrete group equivariance into traditional CNNs through group convolutions. Steerable CNNs (Cohen & Welling, 2017) introduce steerable filters, which provide a more flexible and efficient way to achieve equivariance compared with GCNNs. Subsequently, SFCNNs (Weiler et al., 2018b) and E(2)-equivariant steerable CNNs (Weiler & Cesa, 2019) extend GCNNs and steerable CNNs to continuous group equivariance, while 3D Steerable CNNs (Weiler et al., 2018a) extend these models to 3D volumetric data. On the other hand, some works use partial differential operators (PDOs) to construct equivariant networks (Shen et al., 2020, 2021, 2022; He et al., 2022; Li et al., 2024b, 2025). Furthermore, equivariant self-supervised learning (Wang et al., 2020; Dangovski et al., 2022; Lee et al., 2022; Garrido et al., 2023; Gupta et al., 2024) has also achieved outstanding results. Based on these theoretical frameworks, equivariant networks are widely applied in various fields, such as mathematics (Zhao et al., 2023), physics (Wang et al., 2021; Hu et al., 2025), biochemistry (Bekkers et al., 2018; Winkels & Cohen, 2019; Graham et al., 2020), and others. EMLP (Finzi et al., 2021) embeds matrix group equivariance into MLPs layerwise, which we discuss in detail in Appendix C.
@@ 4.1 Space Structures
where $p_{i,a},q_{i,a},p_{o,a},q_{o,a},c_{i},c_{o},A_{i},A_{o}\in\mathbb{N}$, $p_{i,a}+q_{i,a}>0$, $p_{o,a}+q_{o,a}>0$, and $cT_{0}=\underbrace{T_{0}\oplus T_{0}\oplus\dots\oplus T_{0}}_{c}$.
We have to emphasize that the actual input/output feature does not lie within $U_{i}$$U_{o}$. To align with gated basis functions, we add a gate scalar $T_{0}$ to each non-scalar term $T(p_{i,a},q_{i,a})$$T(p_{o,a},q_{o,a})$ in $U_{i}$$U_{o}$ to obtain the actual input/output space. The mechanism behind this approach will be discussed in detail in Section 4.2. We denote this actual input/output space as the gated input/output space $U_{gi}$$U_{go}$:
@@ 4.2 Gated Basis Functions
The following theorem guarantees the equivariance between the gated input space and the post-activation space (see Appendix D for proof).
@@ 4.3 Equivariant Linear Weights
where $\rho_{go,i}=\rho_{go}\otimes\rho_{i}^{*}$ is the group representation of $U_{go}\otimes U_{i}^{*}$. We use the same method as EMLP (Finzi et al., 2021) to solve for the equivariant basis $Q$ and equivariant projector $P=QQ^{\top}$ of $\mathrm{vec}(W_{b})$. Similar to the transition from Equation 27 to Equation 28 (in Appendix C), we decompose the group representation $\rho_{go,i}(g)$ into discrete and infinitesimal generators to obtain the following constraint:
@@
We summarize the architecture of the EKAN layer in Figure 2 (Right). In this example, the EKAN layer is equivariant with respect to a 2-dimensional matrix group $\widetilde{G}$ (such as the $\mathrm{SO}(2)$ group). The user specifies the input space $U_{i}=T_{0}\oplus T_{1}$ (where we abbreviate $T(p,0)=V^{p}$ as $T_{p}$), which represents a scalar space and a vector space, and specifies the output space $U_{o}=T_{2}$, which represents a matrix space. Then, the gated input space $U_{gi}=T_{0}\oplus T_{1}\oplus T_{0}$ and the gated output space $U_{go}=T_{2}\oplus T_{0}$ each add a gate scalar to the vector space $T_{1}$ and the matrix space $T_{2}$. The basis functions are applied to the gate scalars of $T_{0}$ (itself) and $T_{1}$, which are then multiplied by the original terms to obtain the post-activation space $U_{m}=3U_{i}$. The linear weights $W\in\mathbb{R}^{5\times 9}$ between $U_{m}$ and $U_{go}$ are within the subspace determined by Section 4.3 to ensure equivariance. Similar to KANs (Liu et al., 2024b), EKAN updates grids based on the input activations, which we discuss in detail in Appendix E.
@@ 5 EKAN Architecture
However, the actual input/output features of the EKAN layer stack lie in $U_{gi}$$U_{go}$. Therefore, we need to add extensions to align the network with the dataset. First, the gate scalars of the actual output feature are directly dropped to obtain the final output label of EKAN, which resides in $U_{o}$. Then, we add a lift layer before the first layer to preprocess the raw input feature of EKAN, which is essentially an equivariant linear layer between $U_{i}$ and $U_{gi}$ (see Section 4.3).
@@ 6 Experiments
In this section, we evaluate the performance of EKAN on regression and classification tasks with known symmetries. Compared with general models such as MLPs, KANs, and equivariant architectures like EMLP (Finzi et al., 2021) and CGENN (Ruhe et al., 2023), EKAN achieves lower test loss and higher test accuracy with smaller datasets or fewer parameters. Additionally, for all trained models in this section, we evaluate their equivariant errors in Appendix F.
@@ 6.1 Particle Scattering
In electron-muon scattering, we can observe the four-momenta of the incoming electron, incoming muon, outgoing electron, and outgoing muon, denoted as $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\in\mathbb{R}^{4}$, respectively. We aim to predict the matrix element, which is proportional to the cross-section (Finzi et al., 2021): $|\mathcal{M}|^{2}\propto[p^{\mu}\tilde{p}^{\nu}-(p^{\alpha}\tilde{p}_{\alpha}-p^{\alpha}p_{\alpha})g^{\mu\nu}][q_{\mu}\tilde{q}_{\nu}-(q^{\alpha}\tilde{q}_{\alpha}-q^{\alpha}q_{\alpha})g_{\mu\nu}]$. According to Einstein’s summation convention, in a monomial, if an index appears once as a superscript and once as a subscript, it indicates summation over that index. The metric tensor is given by $g_{\mu\nu}=g^{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$, and $a_{\mu}=g_{\mu\nu}a^{\nu}=(a^{0},-a^{1},-a^{2},-a^{3})$. The matrix element is invariant under Lorentz transformations. In other words, this task exhibits $\mathrm{O}(1,3)$ invariance (see Section A.2 for more details), with the feature space $U_{i}=4T_{1}$ and the label space $U_{o}=T_{0}$. The data generation process for particle scattering is entirely consistent with that in EMLP (Finzi et al., 2021).
We embed the group $\mathrm{O}(1,3)$ and its subgroups $\mathrm{SO}^{+}(1,3)$ and $\mathrm{SO}(1,3)$ equivariance into EKAN. Models are evaluated on synthetic datasets with different training set sizes, which are generated by sampling $q^{\mu},p^{\mu},\tilde{q}^{\mu},\tilde{p}^{\mu}\sim\mathcal{N}(0,\frac{1}{4^{2}})$. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$. Although the lift layer increases the parameter overhead, we set the width of the middle layer in EKAN to $n_{1}=1000$ (shape as $[16,1000,1]$, and the software will automatically calculate the appropriate feature space structure based on the user-specified dimension), and set the width of the middle layer in KAN to $n_{1}=3840$ (shape as $[16,3840,1]$) to keep the parameter count similar. Both EMLP and MLP have the depth of $L=4$ and the middle layer width of $n_{1}=n_{2}=n_{3}=384$ (shape as $[16,384,384,384,1]$). In these settings, EKAN (435435k) has fewer parameters than EMLP (450450k) and KAN (461461k). We provide more implementation details in Section G.1.
@@ 6.2 Three-Body Problem
Table 3: Test accuracy ($\%$) of different models on the top quark tagging dataset ($n_{comp}=3$) with different training set sizes. We present the results in the format of mean $\pm$ std.
@@
EMLP-$\mathrm{SO}^{+}(1,3)$ $65.48\pm 1.21$ $72.59\pm 0.84$ $74.40\pm 0.26$ $6.34\pm 0.14 77$$7.10\pm 0.02$ 133133K
[…]
KAN $49.89\pm 0.39$ $49.91\pm 0.43$ $49.89\pm 0.37$ $50.00\pm 0.02$ $49.84\pm 0.25$ 3535K
@@
CGENN $62.63\pm 2.24$ $68.74\pm 0.77$ $70.29\pm 1.29$ $5.10\pm 0.47 77$$7.05\pm 0.03$ 8585K
EKAN-$\mathrm{SO}^{+}(1,3)$ (Ours) $\mathbf{71.92\pm 0.88}$ $\mathbf{73.98\pm 0.39}$ $\mathbf{76.15\pm 0.11}$ $\mathbf{76.69\pm 0.08}$ $76.93\pm 0.02$ 3434K
@@
The study of the three-body problem on a plane (Greydanus et al., 2019) focuses on the motion of three particles, with their center of mass at the origin, under the influence of gravity. Their trajectories are chaotic and cannot be described by an analytical solution. Specifically, we observe the motion states of three particles over the past four time steps, denoted as $\{q_{i1},p_{i1},q_{i2},p_{i2},q_{i3},p_{i3}\}_{i=t-4}^{t-1}$, and predict their motion states at time $t$, denoted as $\{q_{t1},p_{t1},q_{t2},p_{t2},q_{t3},p_{t3}\}$. Here, $q_{ij}\in\mathbb{R}^{2}$ and $p_{ij}\in\mathbb{R}^{2}$ indicate the position and momentum coordinates of the $j$-th particle at time $i$, respectively. The dataset contains 3030k training samples and 3030k test samples. When the input motion states are simultaneously rotated by a certain angle or reflected along a specific axis, the output motion states should undergo the same transformation. Therefore, this task has $\mathrm{O}(2)$ equivariance (see Section A.1 for more details), with the feature space $U_{i}=4\times 6T_{1}=24T_{1}$ and the label space $U_{o}=6T_{1}$. The dataset for the three-body problem comes from HNN (Greydanus et al., 2019), and we note that they predict the motion trajectories of three particles, differing from the setup in CGENN (Ruhe et al., 2023), which addresses an $N$-body problem involving five particles.
We embed the group $\mathrm{O}(2)$ and its subgroup $\mathrm{SO}(2)$ equivariance into EKAN and EMLP. Both EKAN and KAN have the depth of $L=2$, the spline order of $k=3$, and grid intervals of $G=3$, while both EMLP and MLP have the depth of $L=4$. The number of parameters is controlled by adjusting the middle layer width $N$ for comparison (the shape of EKAN and KAN is $[48,N,12]$, while the shape of EMLP and MLP is $[48,N,N,N,12]$). More implementation details can be found in Section G.2.